OpenStudy (anonymous):

use the change base formula to evaluate the logarithm. log v7 12 (small seven then base 12

OpenStudy (anonymous):

\[\log_7(12)=\frac{\log(12)}{\log(7)}\]

OpenStudy (anonymous):

btw the "base" is 7, not 12 because \[\log_7(x)=y\] means \[7^y=x\]

myininaya (myininaya):

what is v?

OpenStudy (anonymous):

and if you like you can also use \[\log_7(12)=\frac{\ln(12)}{\ln(7)}\]

OpenStudy (anonymous):

v is a little arrow pointing down i bet